# Reasoning about the Unseen for Efficient Outdoor Object Navigation

### v1

@@ Abstract
Robots should exist anywhere humans do: indoors, outdoors, and even unmapped environments. In contrast, the focus of recent advancements in Object Goal Navigation (OGN)[1, 2, 3] has targeted navigating in indoor environments by leveraging spatial and semantic cues that do not generalize outdoors. While these contributions provide valuable insights into indoor scenarios, the broader spectrum of real-world robotic applications often extends to outdoor settings. As we transition to the vast and complex terrains of outdoor environments, new challenges emerge. Unlike the structured layouts found indoors, outdoor environments lack clear spatial delineations and are riddled with inherent semantic ambiguities. Despite this, humans navigate with ease because we can reason about the unseen. We introduce a new task OUTDOOR, a new mechanism for Large Language Models (LLMs) to accurately hallucinate possible futures, and a new computationally aware success metric for pushing research forward in this more complex domain. Additionally, we show impressive results on both a simulated drone and physical quadruped in outdoor environments. Our agent has no premapping and our formalism outperforms naive LLM-based approaches.
Figure 1: Direct application of Language Models in embodied agents navigating outdoor environments suffers from short-sightedness and limited environment comprehension. Our approach augments the LLM by enabling it to expand imaginary nodes in space, enhancing feasibility for outdoor navigation.
@@ I Introduction
Advancements in Object Goal Navigation (OGN) [1, 2, 3] have enhanced the proficiency of robotic agents in navigating indoor environments by leveraging spatial and semantic cues. Agents that can guide humans (e.g. the visually impaired [4]) are an important enabling technology, but need to move beyond restricted indoor spaces to the full richness of outdoor navigation. Outdoor environments are substantially larger than handled by current semantic mapping approaches [5, 6], have complex terrains [7], and, crucially, lack clear semantic delineations. Not only is sensing simplified indoors, but so is reasoning as rooms are easily distinguished and semantically categorized. Outdoor environments still have semantic distinctions but visually identical spaces might be a soccer field, a picnic area, or the pit of an outdoor orchestra depending on the time of day. Additionally, outdoor navigational tasks typically demand that robotic agents engage in roles with more granular goal specifications. For instance, in the context of search and rescue operations, the objective is not merely to navigate to a general category of ‘people’ but to pinpoint casualties potentially trapped under a car.
Recently, Large Language Models (LLMs) [8, 9, 10] trained on expansive internet datasets are serving as adaptable policies in embodied platforms, making them proficient in addressing a wider range of tasks [11, 12, 13]. The existing work has primarily focused on high-level task-planning with predefined skills in constrained environments. Despite this, we have seem very promising skill demonstrations in indoor object-scenarios made possible by these models [14, 15, 16]. While some emergent behavior has been identified, the language and vision communities have begun harnessing these LLMs for their reasoning capabilities due to the vast world knowledge and models stored in their parameters.
@@ III-A Decision Making and Planning for LLM
Vanilla implementations of Large Language Models (LLMs) often fall short in decision-making and planning capabilities. To address this, several strategies have been developed. The linear reasoning approach, “Chain of Thoughts,” enhances structured problem-solving [17], and tree-based strategies, “Tree of Thoughts,” bring forth search-guided reasoning capabilities [18]. To improve performance search algorithms have also been integrated [19].
External planning methods have emerged [20, 21, 22, 23, 24, 25, 26] as methods to leverage techniques such as Monte Carlo Tree Search (MCTS) to enhance the reasoning capacities of LLMs. While they show promise in fields like mathematics, code generation, and high-level task planning, their application in low-level path planning for robotics remains limited. A primary reason is the complexity of mapping LLM outputs to the intricate action spaces of robots, making tasks requiring detailed sequences of movements a challenge. In this context, our approach stands out by using waypoints as a natural interface for low-level path planning. This not only bridges the gap between LLM reasoning and robot actions but also ensures that our method operates in a parallel manner, offering both effectiveness and efficiency.
@@ III-B LLM as embodied agents for navigation
The use of language for guiding embodied agents has a long lineage. Language only models like BERT [9] can be used as scoring functions between language instructions and path to help embodied agent navigate [27]. Performance on such tasks have scaled with larger models (e.g. GPT-4) which have greater aptitudes for common sense reasoning and its comprehension of world structures  [15, 14, 16, 12, 28]. Shah et al  [12] uses LLMs as a parser to extract landmarks as sub-goal nodes for robots to navigate on a graph, Chen et al [14] use LLM as an evaluator to re-weight the waypoints generated by the frontier-based method  [29]. Zhou et al  [15] took another approach, they used several hand-designed constraints via Probabilistic Soft Logic programming language to choose the best frontiers to explore. NavGPT [16] utilized the synergizing prompt methods such as ReAct [30] with a discrete action space for LLM to navigate.
Figure 3: Overview: The agent captures $N$ RGB images (potential frontiers). Each image is processed through a Vision Language Model (VLM) to generate a textual caption. Subsequent Rapidly-exploring Random Trees (RRT) aid the agent in envisioning possible future scenarios for each frontier. The results, combined with GPS coordinates, populate a frontier buffer. The most promising frontier is identified, and a local planner guides the agent to its location.
Existing approaches to navigation predominantly rely on idealized indoor scene graphs, often assuming structured environments where, for example, a refrigerator is necessarily located in the kitchen or a fireplace in the living room  [15, 14, 12]. Alternatively, some methods leverage Google Street View data for navigation tasks  [31]. However, these approaches fall short in capturing the nuanced complexities and granularities inherent to real-world scenarios, such as search and rescue operations or advanced domestic robotics tasks in semantically rich environments such as airports or campus buildings. In genuine outdoor settings, spatial semantics may be ambiguous or lack well-defined boundaries. Consequently, an intelligent agent must be capable of strategically predict information in space to effectively navigate and reason within these more complex contexts.
@@ IV-A Graph the unknown
Historically, methods for object goal navigation and VLN relied on near perfect depth information derived from simulations to generate dense geometric occupancy maps, which subsequently informed the expansion of explorable frontiers [14, 15, 16, 32]. However, these assumptions falter in real-world outdoor settings, especially without the aid of high-end depth cameras or LiDARs. Addressing this limitation, our approach introduces an adaptive topological graph to introduce frontiers. This combination not only mitigates the need for perfect depth information but also enhances the agent’s navigational capabilities.
@@ IV-B Reasoning about the uncertainty
In the context of planning within intricate outdoor environments, direct determinations by the LLM-based method solely on localized information can lead to suboptimal behaviors, such as inconsistency and short-sightedness. These behaviors arise not only from the inherent tendency of LLMs to hallucinate [33], but also the non-delineated nature of the outdoor environment. Such properties become particularly concerning when placing heavy reliance on singular output generations from LLMs [34].
To fortify against these vulnerabilities and enhance decision robustness, we elected to incorporate the future information using expanding RRT strategy. By projecting multiple forward-looking imaginary branches through iterative queries of LLM_Visionary and LLM_Evaluator (Figure 3), mitigating the risks associated with single query outputs. Our choice of RRT was informed by its intrinsic properties, in contrast of the sequential sampling process used in MCTS[22][20], RRT provides a parallelizable framework to allow us score and expand the imaginary nodes all at the same time.
@@ IV-C Perceiving
The agent captures $N$ images during each exploration steps. These images are processed by a Vision Language Model (VLM). For the purposes of this study, we employed Kosmos-2 [35], a VLM fine-tuned with spatially-grounded image-text data. This model offers the distinct advantage of providing detailed object-level descriptions of scenes. Importantly, it is promptable, so we prompt it to describe not only the objects in the scene, but also the spatial relationship between objects as well as the backgrounds.
@@ V Experiments
TABLE I: Baseline Comparison for Different Task Levels in simulation (AirSim [36]
Comprehensive evaluations are conducted across multiple platforms: 1. The AirSim [36] simulation environment: A photo-realistic outdoor simulation setting that consists of different semantic distinguishable areas in Downtown West environment. 2. A real-world robotic platform: Unitree Go1, equiped with a USB camera, high resolution RTK-GPS module, and Inertial measurement unit.
@@ V-B Baselines
We benchmarked our approach against two baseline implementations: The first, LLM-as-Eval [14], utilizes the LLM as an evaluator for re-scoring expanded frontiers [29]. The second, LLM-MCTS [20], employs Monte Carlo Tree Search techniques for trajectory expansion at 10 iterations. Both baselines are reproductions of existing methods using our graph maps, devoid of depth information. This choice is informed by real-world practicalities: without advanced depth cameras, standard devices such as the Real Sense D435 exhibit subpar performance in outdoor environments.
@@ V-C Result
Table I presents a comprehensive evaluation of our methods against the baselines on four metrics: Success Rate (SR), Oracle Success Rate (OSR), Success Weighted by Path Length (SPL) [1], and our newly introduced metric, Computationally Adjusted Success Rate (CASR). It shows:

### v2

@@ Abstract
Robots should exist anywhere humans do: indoors, outdoors, and even unmapped environments. In contrast, the focus of recent advancements in Object Goal Navigation (OGN)[anderson2018objectnav, chaplot2020object, majumdar2022zson] has targeted navigating in indoor environments by leveraging spatial and semantic cues that do not generalize outdoors. While these contributions provide valuable insights into indoor scenarios, the broader spectrum of real-world robotic applications often extends to outdoor settings. As we transition to the vast and complex terrains of outdoor environments, new challenges emerge. Unlike the structured layouts found indoors, outdoor environments lack clear spatial delineations and are riddled with inherent semantic ambiguities. Despite this, humans navigate with ease because we can reason about the unseen. We introduce a new task OUTDOOR, a new mechanism for Large Language Models (LLMs) to accurately hallucinate possible futures, and a new computationally aware success metric for pushing research forward in this more complex domain. Additionally, we show impressive results on both a simulated drone and physical quadruped in outdoor environments. Our agent has no premapping and our formalism outperforms naive LLM-based approaches.
@@
figureDirect application of Language Models in embodied agents navigating outdoor environments suffers from short-sightedness and limited environment comprehension. Our approach augments the LLM by enabling it to expand imaginary nodes in space, enhancing feasibility for outdoor navigation.
1Carnegie Mellon University, Pittsburgh, PA USA. {quantinx, tianyiz4, kedix, mkj, ybisk}@andrew.cmu.edu2https://github.com/quantingxie/ReasonedExplorer
@@ I Introduction
Advancements in Object Goal Navigation (OGN) [anderson2018objectnav, chaplot2020object, majumdar2022zson] have enhanced the proficiency of robotic agents in navigating indoor environments by leveraging spatial and semantic cues. Agents that can guide humans (e.g. the visually impaired [blind-indoors]) are an important enabling technology, but need to move beyond restricted indoor spaces to the full richness of outdoor navigation. Outdoor environments are substantially larger than handled by current semantic mapping approaches [chaplot2020learning, Min2022], have complex terrains [self-supervised-outdoor], and, crucially, lack clear semantic delineations. Not only is sensing simplified indoors, but so is reasoning as rooms are easily distinguished and semantically categorized. Outdoor environments still have semantic distinctions but visually identical spaces might be a soccer field, a picnic area, or the pit of an outdoor orchestra depending on the time of day. Additionally, outdoor navigational tasks typically demand that robotic agents engage in roles with more granular goal specifications. For instance, in the context of search and rescue operations, the objective is not merely to navigate to a general category of ‘people’ but to pinpoint casualties potentially trapped under a car.
Recently, Large Language Models (LLMs) [gpt, devlin2018bert, openai2023gpt4] trained on expansive internet datasets are serving as adaptable policies in embodied platforms, making them proficient in addressing a wider range of tasks [saycan2022arxiv, shah2023lmnav, brohan2022rt]. The existing work has primarily focused on high-level task-planning with predefined skills in constrained environments. Despite this, we have seem very promising skill demonstrations in indoor object-scenarios made possible by these models [chen2023train, zhou2023esc, zhou2023navgpt]. While some emergent behavior has been identified, the language and vision communities have begun harnessing these LLMs for their reasoning capabilities due to the vast world knowledge and models stored in their parameters.
@@ III-A Decision Making and Planning for LLM
Vanilla implementations of Large Language Models (LLMs) often fall short in decision-making and planning capabilities. To address this, several strategies have been developed. The linear reasoning approach, “Chain of Thoughts,” enhances structured problem-solving [wei2023chainofthought], and tree-based strategies, “Tree of Thoughts,” bring forth search-guided reasoning capabilities [yao2023tree]. To improve performance search algorithms have also been integrated [xie2023decomposition].
External planning methods have emerged [hao2023reasoning, zhao2023large, zhang2023planning, wang2023planandsolve, wang2023describe, huang2022inner, yao2023react] as methods to leverage techniques such as Monte Carlo Tree Search (MCTS) to enhance the reasoning capacities of LLMs. While they show promise in fields like mathematics, code generation, and high-level task planning, their application in low-level path planning for robotics remains limited. A primary reason is the complexity of mapping LLM outputs to the intricate action spaces of robots, making tasks requiring detailed sequences of movements a challenge. In this context, our approach stands out by using waypoints as a natural interface for low-level path planning. This not only bridges the gap between LLM reasoning and robot actions but also ensures that our method operates in a parallel manner, offering both effectiveness and efficiency.
@@ III-B LLM as embodied agents for navigation
The use of language for guiding embodied agents has a long lineage. Language only models like BERT [devlin2018bert] can be used as scoring functions between language instructions and path to help embodied agent navigate [majumdar2020vlnbert]. Performance on such tasks have scaled with larger models (e.g. GPT-4) which have greater aptitudes for common sense reasoning and its comprehension of world structures  [zhou2023esc, chen2023train, zhou2023navgpt, shah2023lmnav, dorbala2023catshapedmug]. Shah et al  [shah2023lmnav] uses LLMs as a parser to extract landmarks as sub-goal nodes for robots to navigate on a graph, Chen et al [chen2023train] use LLM as an evaluator to re-weight the waypoints generated by the frontier-based method  [1997frontier]. Zhou et al  [zhou2023esc] took another approach, they used several hand-designed constraints via Probabilistic Soft Logic programming language to choose the best frontiers to explore. NavGPT [zhou2023navgpt] utilized the synergizing prompt methods such as ReAct [yao2022react] with a discrete action space for LLM to navigate.
[…]
figureOverview: The agent captures $N$ RGB images (potential frontiers). Each image is processed through a Vision Language Model (VLM) to generate a textual caption. Subsequent Rapidly-exploring Random Trees (RRT) aid the agent in envisioning possible future scenarios for each frontier. The results, combined with GPS coordinates, populate a frontier buffer. The most promising frontier is identified, and a local planner guides the agent to its location.
Existing approaches to navigation predominantly rely on idealized indoor scene graphs, often assuming structured environments where, for example, a refrigerator is necessarily located in the kitchen or a fireplace in the living room  [zhou2023esc, chen2023train, shah2023lmnav]. Alternatively, some methods leverage Google Street View data for navigation tasks  [schumann2023velma]. However, these approaches fall short in capturing the nuanced complexities and granularities inherent to real-world scenarios, such as search and rescue operations or advanced domestic robotics tasks in semantically rich environments such as airports or campus buildings. In genuine outdoor settings, spatial semantics may be ambiguous or lack well-defined boundaries. Consequently, an intelligent agent must be capable of strategically predict information in space to effectively navigate and reason within these more complex contexts.
@@ IV-A Graph the unknown
Historically, methods for object goal navigation and VLN relied on near perfect depth information derived from simulations to generate dense geometric occupancy maps, which subsequently informed the expansion of explorable frontiers [chen2023train, zhou2023esc, zhou2023navgpt, pmlr-v155-anderson21a]. However, these assumptions falter in real-world outdoor settings, especially without the aid of high-end depth cameras or LiDARs. Addressing this limitation, our approach introduces an adaptive topological graph to introduce frontiers. This combination not only mitigates the need for perfect depth information but also enhances the agent’s navigational capabilities.
@@ IV-B Reasoning about the uncertainty
In the context of planning within intricate outdoor environments, direct determinations by the LLM-based method solely on localized information can lead to suboptimal behaviors, such as inconsistency and short-sightedness. These behaviors arise not only from the inherent tendency of LLMs to hallucinate [McKenna2023SourcesOH], but also the non-delineated nature of the outdoor environment. Such properties become particularly concerning when placing heavy reliance on singular output generations from LLMs [stochastic].
To fortify against these vulnerabilities and enhance decision robustness, we elected to incorporate the future information using expanding RRT strategy. By projecting multiple forward-looking imaginary branches through iterative queries of LLM_Visionary and LLM_Evaluator (Figure III-B), mitigating the risks associated with single query outputs. Our choice of RRT was informed by its intrinsic properties, in contrast of the sequential sampling process used in MCTS[zhang2023planning][hao2023reasoning], RRT provides a parallelizable framework to allow us score and expand the imaginary nodes all at the same time.
@@ IV-C Perceiving
The agent captures $N$ images during each exploration steps. These images are processed by a Vision Language Model (VLM). For the purposes of this study, we employed Kosmos-2 [kosmos-2], a VLM fine-tuned with spatially-grounded image-text data. This model offers the distinct advantage of providing detailed object-level descriptions of scenes. Importantly, it is promptable, so we prompt it to describe not only the objects in the scene, but also the spatial relationship between objects as well as the backgrounds.
@@ V Experiments
TABLE I: Baseline Comparison for Different Task Levels in simulation (AirSim [shah2017airsim]
Comprehensive evaluations are conducted across multiple platforms: 1. The AirSim [shah2017airsim] simulation environment: A photo-realistic outdoor simulation setting that consists of different semantic distinguishable areas in Downtown West environment. 2. A real-world robotic platform: Unitree Go1, equiped with a USB camera, high resolution RTK-GPS module, and Inertial measurement unit.
@@ V-B Baselines
We benchmarked our approach against two baseline implementations: The first, LLM-as-Eval [chen2023train], utilizes the LLM as an evaluator for re-scoring expanded frontiers [1997frontier]. The second, LLM-MCTS [hao2023reasoning], employs Monte Carlo Tree Search techniques for trajectory expansion at 10 iterations. Both baselines are reproductions of existing methods using our graph maps, devoid of depth information. This choice is informed by real-world practicalities: without advanced depth cameras, standard devices such as the Real Sense D435 exhibit subpar performance in outdoor environments.
@@ V-C Result
Table I presents a comprehensive evaluation of our methods against the baselines on four metrics: Success Rate (SR), Oracle Success Rate (OSR), Success Weighted by Path Length (SPL) [anderson2018objectnav], and our newly introduced metric, Computationally Adjusted Success Rate (CASR). It shows: